\begin{thebibliography}{38}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Aluru et~al.(2020)Aluru, Mathew, Saha, and Mukherjee}]{aluru2020deep}
Sai~Saket Aluru, Binny Mathew, Punyajoy Saha, and Animesh Mukherjee. 2020.
\newblock Deep learning models for multilingual hate speech detection.
\newblock \emph{arXiv preprint arXiv:2004.06465}.

\bibitem[{Basile et~al.(2019)Basile, Bosco, Fersini, Debora, Patti, Pardo,
  Rosso, Sanguinetti et~al.}]{basile2019semeval}
Valerio Basile, Cristina Bosco, Elisabetta Fersini, Nozza Debora, Viviana
  Patti, Francisco Manuel~Rangel Pardo, Paolo Rosso, Manuela Sanguinetti,
  et~al. 2019.
\newblock Semeval-2019 task 5: Multilingual detection of hate speech against
  immigrants and women in twitter.
\newblock In \emph{13th International Workshop on Semantic Evaluation}, pages
  54--63. Association for Computational Linguistics.

\bibitem[{Bosco et~al.(2018)Bosco, Felice, Poletto, Sanguinetti, and
  Maurizio}]{bosco2018overview}
Cristina Bosco, Dell'Orletta Felice, Fabio Poletto, Manuela Sanguinetti, and
  Tesconi Maurizio. 2018.
\newblock Overview of the evalita 2018 hate speech detection task.
\newblock In \emph{EVALITA 2018-Sixth Evaluation Campaign of Natural Language
  Processing and Speech Tools for Italian}, volume 2263, pages 1--9. CEUR.

\bibitem[{Castelle(2018)}]{castelle2018linguistic}
Michael Castelle. 2018.
\newblock The linguistic ideologies of deep abusive language classification.
\newblock In \emph{Proceedings of the 2nd Workshop on Abusive Language Online
  (ALW2)}, pages 160--170.

\bibitem[{Chakrabarty et~al.(2019)Chakrabarty, Gupta, and
  Muresan}]{chakrabarty2019pay}
Tuhin Chakrabarty, Kilol Gupta, and Smaranda Muresan. 2019.
\newblock Pay “attention” to your context when classifying abusive
  language.
\newblock In \emph{Proceedings of the Third Workshop on Abusive Language
  Online}, pages 70--79.

\bibitem[{Da~San~Martino et~al.(2019)Da~San~Martino, Yu, Barr{\'o}n-Cedeno,
  Petrov, and Nakov}]{da2019fine}
Giovanni Da~San~Martino, Seunghak Yu, Alberto Barr{\'o}n-Cedeno, Rostislav
  Petrov, and Preslav Nakov. 2019.
\newblock Fine-grained analysis of propaganda in news article.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 5640--5650.

\bibitem[{Devlin et~al.(2018)Devlin, Chang, Lee, and
  Toutanova}]{DBLP:journals/corr/abs-1810-04805}
Jacob Devlin, Ming{-}Wei Chang, Kenton Lee, and Kristina Toutanova. 2018.
\newblock \href {http://arxiv.org/abs/1810.04805} {{BERT:} pre-training of deep
  bidirectional transformers for language understanding}.
\newblock \emph{CoRR}, abs/1810.04805.

\bibitem[{Fortuna and Nunes(2018)}]{fortuna2018survey}
Paula Fortuna and S{\'e}rgio Nunes. 2018.
\newblock A survey on automatic detection of hate speech in text.
\newblock \emph{ACM Computing Surveys (CSUR)}, 51(4):1--30.

\bibitem[{Fortuna et~al.(2020)Fortuna, Soler, and Wanner}]{fortuna2020toxic}
Paula Fortuna, Juan Soler, and Leo Wanner. 2020.
\newblock Toxic, hateful, offensive or abusive? what are we really classifying?
  an empirical analysis of hate speech datasets.
\newblock In \emph{Proceedings of the 12th Language Resources and Evaluation
  Conference}, pages 6786--6794.

\bibitem[{Harris(1954)}]{harris1954distributional}
Zellig~S Harris. 1954.
\newblock Distributional structure.
\newblock \emph{Word}, 10(2-3):146--162.

\bibitem[{Ibrohim and Budi(2018)}]{ibrohim2018dataset}
Muhammad~Okky Ibrohim and Indra Budi. 2018.
\newblock A dataset and preliminaries study for abusive language detection in
  indonesian social media.
\newblock \emph{Procedia Computer Science}, 135:222--229.

\bibitem[{Jiao et~al.(2020)Jiao, Yin, Shang, Jiang, Chen, Li, Wang, and
  Liu}]{jiao2020tinybert}
Xiaoqi Jiao, Yichun Yin, Lifeng Shang, Xin Jiang, Xiao Chen, Linlin Li, Fang
  Wang, and Qun Liu. 2020.
\newblock Tinybert: Distilling bert for natural language understanding.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing: Findings}, pages 4163--4174.

\bibitem[{Karan and {\v{S}}najder(2018)}]{karan2018cross}
Mladen Karan and Jan {\v{S}}najder. 2018.
\newblock Cross-domain detection of abusive language online.
\newblock In \emph{Proceedings of the 2nd workshop on abusive language online
  (ALW2)}, pages 132--137.

\bibitem[{Komalova et~al.(2021)Komalova, Goloshchapova, Motovskikh, Epifanov,
  Morozov, and Glazkova}]{komalova2021mca}
Liliya Komalova, Tatiana Goloshchapova, Leonid Motovskikh, Rostislav Epifanov,
  Dmitry Morozov, and Anna Glazkova. 2021.
\newblock \href {https://doi.org/10.17632/fktgy52645.1} {Mca workshop - toxic
  comments}.
\newblock Mendeley Data, V1.

\bibitem[{Lee et~al.(2018)Lee, Yoon, and Jung}]{lee2018comparative}
Younghun Lee, Seunghyun Yoon, and Kyomin Jung. 2018.
\newblock Comparative studies of detecting abusive language on twitter.
\newblock In \emph{Proceedings of the 2nd Workshop on Abusive Language Online
  (ALW2)}, pages 101--106.

\bibitem[{Leite et~al.(2020)Leite, Silva, Bontcheva, and
  Scarton}]{leite2020toxic}
JA~Leite, DF~Silva, K~Bontcheva, and C~Scarton. 2020.
\newblock Toxic language detection in social media for brazilian portuguese:
  new dataset and multilingual analysis.
\newblock In \emph{Proceedings of the 1st Conference of the Asia-Pacific
  Chapter of the Association for Computational Linguistics and the 10th
  International Joint Conference on Natural Language Processing}, pages
  914--924. Association for Computational Linguistics (ACL).

\bibitem[{Liu et~al.(2019{\natexlab{a}})Liu, Li, and Zou}]{liu2019nuli}
Ping Liu, Wen Li, and Liang Zou. 2019{\natexlab{a}}.
\newblock Nuli at semeval-2019 task 6: Transfer learning for offensive language
  detection using bidirectional transformers.
\newblock In \emph{Proceedings of the 13th international workshop on semantic
  evaluation}, pages 87--91.

\bibitem[{Liu et~al.(2019{\natexlab{b}})Liu, Ott, Goyal, Du, Joshi, Chen, Levy,
  Lewis, Zettlemoyer, and Stoyanov}]{liu2019roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019{\natexlab{b}}.
\newblock Roberta: A robustly optimized bert pretraining approach.
\newblock \emph{arXiv preprint arXiv:1907.11692}.

\bibitem[{Loshchilov and Hutter(2017)}]{loshchilov2017decoupled}
Ilya Loshchilov and Frank Hutter. 2017.
\newblock Decoupled weight decay regularization.
\newblock \emph{arXiv preprint arXiv:1711.05101}.

\bibitem[{Mikolov et~al.(2013)Mikolov, Chen, Corrado, and
  Dean}]{mikolov2013efficient}
Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013.
\newblock Efficient estimation of word representations in vector space.
\newblock \emph{arXiv preprint arXiv:1301.3781}.

\bibitem[{Mozafari et~al.(2019)Mozafari, Farahbakhsh, and
  Crespi}]{mozafari2019bert}
Marzieh Mozafari, Reza Farahbakhsh, and Noel Crespi. 2019.
\newblock A bert-based transfer learning approach for hate speech detection in
  online social media.
\newblock In \emph{International Conference on Complex Networks and Their
  Applications}, pages 928--940. Springer.

\bibitem[{Mubarak et~al.(2020)Mubarak, Darwish, Magdy, Elsayed, and
  Al-Khalifa}]{mubarak2020overview}
Hamdy Mubarak, Kareem Darwish, Walid Magdy, Tamer Elsayed, and Hend Al-Khalifa.
  2020.
\newblock Overview of osact4 arabic offensive language detection shared task.
\newblock In \emph{Proceedings of the 4th Workshop on Open-Source Arabic
  Corpora and Processing Tools, with a Shared Task on Offensive Language
  Detection}, pages 48--52.

\bibitem[{Nikolov and Radivchev(2019)}]{nikolov2019nikolov}
Alex Nikolov and Victor Radivchev. 2019.
\newblock Nikolov-radivchev at semeval-2019 task 6: Offensive tweet
  classification with bert and ensembles.
\newblock In \emph{Proceedings of the 13th International Workshop on Semantic
  Evaluation}, pages 691--695.

\bibitem[{Park and Fung(2017)}]{park2017one}
Ji~Ho Park and Pascale Fung. 2017.
\newblock One-step and two-step classification for abusive language detection
  on twitter.
\newblock In \emph{Proceedings of the First Workshop on Abusive Language
  Online}, pages 41--45.

\bibitem[{Pavlopoulos et~al.(2021)Pavlopoulos, Laugier, Sorensen, and
  Androutsopoulos}]{pav2020semeval}
John Pavlopoulos, Léo Laugier, Jeffrey Sorensen, and Ion Androutsopoulos.
  2021.
\newblock Semeval-2021 task 5: Toxic spans detection (to appear).
\newblock In \emph{Proceedings of the 15th International Workshop on Semantic
  Evaluation}.

\bibitem[{Pavlopoulos et~al.(2017{\natexlab{a}})Pavlopoulos, Malakasiotis, and
  Androutsopoulos}]{pavlopoulos2017deep}
John Pavlopoulos, Prodromos Malakasiotis, and Ion Androutsopoulos.
  2017{\natexlab{a}}.
\newblock Deep learning for user comment moderation.
\newblock In \emph{Proceedings of the First Workshop on Abusive Language
  Online}, pages 25--35.

\bibitem[{Pavlopoulos et~al.(2017{\natexlab{b}})Pavlopoulos, Malakasiotis, and
  Androutsopoulos}]{pavlopoulos2017deeper}
John Pavlopoulos, Prodromos Malakasiotis, and Ion Androutsopoulos.
  2017{\natexlab{b}}.
\newblock Deeper attention to abusive user content moderation.
\newblock In \emph{Proceedings of the 2017 conference on empirical methods in
  natural language processing}, pages 1125--1135.

\bibitem[{Pitenis et~al.(2020)Pitenis, Zampieri, and
  Ranasinghe}]{pitenis2020offensive}
Zesis Pitenis, Marcos Zampieri, and Tharindu Ranasinghe. 2020.
\newblock Offensive language identification in greek.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 5113--5119.

\bibitem[{Risch et~al.(2019)Risch, Stoll, Ziegele, and
  Krestel}]{risch2019hpidedis}
Julian Risch, Anke Stoll, Marc Ziegele, and Ralf Krestel. 2019.
\newblock hpidedis at germeval 2019: Offensive language identification using a
  german bert model.
\newblock In \emph{KONVENS}.

\bibitem[{Rosenthal et~al.(2020)Rosenthal, Atanasova, Karadzhov, Zampieri, and
  Nakov}]{rosenthal2020large}
Sara Rosenthal, Pepa Atanasova, Georgi Karadzhov, Marcos Zampieri, and Preslav
  Nakov. 2020.
\newblock A large-scale semi-supervised dataset for offensive language
  identification.
\newblock \emph{arXiv preprint arXiv:2004.14454}.

\bibitem[{Sanh et~al.(2019)Sanh, Debut, Chaumond, and
  Wolf}]{sanh2019distilbert}
Victor Sanh, Lysandre Debut, Julien Chaumond, and Thomas Wolf. 2019.
\newblock Distilbert, a distilled version of bert: smaller, faster, cheaper and
  lighter.
\newblock \emph{arXiv preprint arXiv:1910.01108}.

\bibitem[{Schmidt and Wiegand(2017)}]{schmidt2017survey}
Anna Schmidt and Michael Wiegand. 2017.
\newblock A survey on hate speech detection using natural language processing.
\newblock In \emph{Proceedings of the fifth international workshop on natural
  language processing for social media}, pages 1--10.

\bibitem[{Sun et~al.(2019)Sun, Qiu, Xu, and
  Huang}]{DBLP:journals/corr/abs-1905-05583}
Chi Sun, Xipeng Qiu, Yige Xu, and Xuanjing Huang. 2019.
\newblock \href {http://arxiv.org/abs/1905.05583} {How to fine-tune {BERT} for
  text classification?}
\newblock \emph{CoRR}, abs/1905.05583.

\bibitem[{Swamy et~al.(2019)Swamy, Jamatia, and
  Gamb{\"a}ck}]{swamy2019studying}
Steve~Durairaj Swamy, Anupam Jamatia, and Bj{\"o}rn Gamb{\"a}ck. 2019.
\newblock Studying generalisability across abusive language detection datasets.
\newblock In \emph{Proceedings of the 23rd Conference on Computational Natural
  Language Learning (CoNLL)}, pages 940--950.

\bibitem[{Waseem et~al.(2017)Waseem, Davidson, Warmsley, and
  Weber}]{waseem2017understanding}
Zeerak Waseem, Thomas Davidson, Dana Warmsley, and Ingmar Weber. 2017.
\newblock Understanding abuse: A typology of abusive language detection
  subtasks.
\newblock In \emph{Proceedings of the First Workshop on Abusive Language
  Online}, pages 78--84.

\bibitem[{Wiegand et~al.(2018)Wiegand, Siegel, and
  Ruppenhofer}]{wiegand2018overview}
Michael Wiegand, Melanie Siegel, and Josef Ruppenhofer. 2018.
\newblock Overview of the germeval 2018 shared task on the identification of
  offensive language.

\bibitem[{Wolf et~al.(2019)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz et~al.}]{wolf2019huggingface}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, R{\'e}mi Louf, Morgan Funtowicz,
  et~al. 2019.
\newblock Huggingface's transformers: State-of-the-art natural language
  processing.
\newblock \emph{arXiv preprint arXiv:1910.03771}.

\bibitem[{Zampieri et~al.(2019)Zampieri, Malmasi, Nakov, Rosenthal, Farra, and
  Kumar}]{zampieri2019semeval}
Marcos Zampieri, Shervin Malmasi, Preslav Nakov, Sara Rosenthal, Noura Farra,
  and Ritesh Kumar. 2019.
\newblock Semeval-2019 task 6: Identifying and categorizing offensive language
  in social media (offenseval).
\newblock \emph{arXiv preprint arXiv:1903.08983}.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{acl2021}
\usepackage{times}
\usepackage{latexsym}
\usepackage{amsmath}

\usepackage{microtype}

\aclfinalcopy 

\title{MIPT-NSU-UTMN at SemEval-2021 Task 5: Ensembling Learning with Pre-trained Language Models for Toxic Spans Detection}

\author{
  Mikhail Kotyushev \\
  Moscow Institute of \\Physics and Technology\\Moscow, Russia \\
  \texttt{mkotyushev@gmail.com} \\\And
  Anna Glazkova \\
  University of \\Tyumen\\Tyumen, Russia \\
  \texttt{a.v.glazkova@utmn.ru} \\\And
  Dmitry Morozov \\
  Novosibirsk State \\University\\Novosibirsk, Russia \\
  \texttt{morozowdm@gmail.com} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
This paper describes our system for SemEval-2021 Task 5 on Toxic Spans Detection. We developed ensemble models using BERT-based neural architectures and post-processing to combine tokens into spans. We evaluated several pre-trained language models using various ensemble techniques for toxic span identification and achieved sizable improvements over our baseline fine-tuned BERT models. Finally, our system obtained a F1-score of 67.55\% on test data.
\end{abstract}

\section{Introduction}

Toxic speech has become a rising issue for social media communities. Abusive content is very diverse and therefore offensive language and toxic speech detection is not a trivial issue. Besides, social media moderation of lengthy comments and posts is often a time-consuming  process. In this regard, the task of detecting toxic spans in social media texts deserves close attention.

This work is based on the participation of our team, named MIPT-NSU-UTMN, in SemEval 2021 Task 5, ``Toxic Spans Detection'' \citep{pav2020semeval}. Organizers of the shared task provided participants with the trial, train, and test sets of English social media comments annotated at the span level indicating the presence or absence of text toxicity. We formulated the task as a token classification problem and investigated several BERT-based models using two-step knowledge transfer. We found that preliminary fine-tuning of the model on data that is close to the target domain improves the quality of the token classification. The source code of our models is available at \url{https://github.com/morozowdmitry/semeval21}.

The paper is organized as follows. A brief review of related work is given in Section 2. The definition of the task has been summarized in Section 3. The proposed methods and experimental settings have been elaborated in Section 4. Section 5 contains the results and error analysis respectively. Section 6 is a conclusion.

\section{Related Work}

Computational approaches to tackle text toxicity have recently gained a lot of interest due to the widespread use of social media. Since moderation is crucial to promoting healthy online discussions, research on toxicity detection has been attracting much attention. Our work is also related to hate speech and abusive language detection \citep{fortuna2020toxic}. 

The toxic speech detection task is usually framed as a supervised learning problem. Moreover, fairly generic features, such as bag of words \citep{harris1954distributional} or word embeddings \citep{mikolov2013efficient}, systematically yield reasonable classification performance \citep{fortuna2018survey,schmidt2017survey}. To better understand the mechanisms of toxic speech detection, some scholars \citep{waseem2017understanding,lee2018comparative,karan2018cross,swamy2019studying} compared different techniques for abusive language analysis. Neural architectures and deep learning methods achieved high results in this domain. Thus, \citet{pavlopoulos2017deep,pavlopoulos2017deeper} explored the possibilities of deep learning and deep attention mechanisms for abusive comment moderation. \citet{park2017one} proposed an approach to performing classification on abusive language based on convolutional neural networks (CNN). \citet{chakrabarty2019pay} used Bidirectional Long-Short Term
Memory network. \citet{castelle2018linguistic} experimented with CNN and Gated Recurrent Units. Some recent studies \citep{mozafari2019bert,risch2019hpidedis,liu2019nuli,nikolov2019nikolov} utilized pre-trained language models such as Bidirectional Encoder Representations from Transformers (BERT) \citep{DBLP:journals/corr/abs-1810-04805} to detect offensive or abusive language.

In recent years, the task of detecting and analyzing abusive, toxic, or offensive language has attracted the attention of more and more researchers. The shared tasks based on carefully curated resources, such as those organized at the SemEval \citep{zampieri2019semeval,basile2019semeval}, GermEval \citep{wiegand2018overview}, EVALITA \citep{bosco2018overview}, and OSACT \citep{mubarak2020overview} events, have significantly contributed to the progress of the field and to the enrichment of linguistic resources. In addition to the corpora collected for these shared tasks, \citet{rosenthal2020large} released a large-scale dataset for offensive language identification. \citet{ibrohim2018dataset,leite2020toxic,pitenis2020offensive,komalova2021mca} presented various datasets for abusive speech detection in non-English languages. Most of these datasets classify whole texts or documents, and do not identify the spans that make a text toxic.

\section{Shared Task}

The task focuses on the evaluation of systems that detect the spans that make a text toxic, when detecting such spans is possible. The goal of the task is to define a sequence of words (character offsets) that attribute to the toxicity of the text, for example:

\begin{itemize}
    \item \textbf{Input.} ``This is a \underline{stupid} example, so thank you for nothing \underline{a!@\#!@}''.
    \item \textbf{Output.} [10,11,12,13,14,15,51,52,53,54,55, 56].
\end{itemize}

The sources of data were various posts (comments) from publicly available datasets. The provided dataset contains 10,629 posts split into training (7939), trial (690), and test (2000) subsets.

Inspired by \citet{da2019fine}, the organizers proposed to employ the F1-score for evaluating the responses of a system participating in the shared task. Let system $A_i$ return a set $S^t_{A_i}$ of character offsets, for parts of the post found to be toxic. Let $G^t$ be the character offsets of the ground truth annotations of $t$. The F1-score of system $A_i$ is calculated with respect to the ground truth $G$ for post $t$ as follows, where $|\cdot|$ denotes set cardinality.

\begin{center}
$\textit{F}^t_1(A_i,G)=\frac{2\cdot P^t(A_i,G)\cdot R^t(A_i,G)}{P^t(A_i,G)+R^t(A_i,G)},$ \newline $\textit{P}^t(A_i,G)=\frac{S^t_{A_i}\cap S^t_G}{S^t_{A_i}},$\newline $\textit{R}^t(A_i,G)=\frac{S^t_{A_i}\cap S^t_G}{S^t_G}.$
\end{center}

The final F1-score is an average $\textit{F}^t_1(A_i,G)$ over all the posts $t$ of an evaluation dataset $T$ to obtain a single score for system $A_i$.

\section{Methodology}
The stated problem was modified from char-level to token-level binary-classification. The proposed solution utilizes a pre-trained language model with a classification head to classify tokens. Different configurations of BERT pre-trained as masked language models were considered as a backbone.

Due to the lack of available token-level labeled public datasets for toxic comment and the relatively small size and sparsity of dataset provided by the competition, the following training pipeline was proposed to enhance knowledge transfer. First, fine-tune pre-trained BERT on a larger-scale task of \textit{toxic comment classification}, using the Jigsaw dataset\footnote{\url{https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge}} from which the competition data were constructed. Second, fine-tune obtained model to solve the actual \textit{toxic tokens classification} problem. The exact training parameters are to be found below.

For the first step:
\begin{itemize}
    \item remove texts occurred in spans dataset from classification dataset to prevent data leakage (so as spans dataset is sampled from classification dataset);
    \item 4 epochs, 200 tokens max length, 64 batch size, 10 gradient accumulation, mixed-precision FP16;
    \item default AdamW \citep{loshchilov2017decoupled} with lr = 4e-5, Layer-wise Decreasing Layer Rate \citep{DBLP:journals/corr/abs-1905-05583} with decay $\eta = 0.95$ and cosine learning rate (LR) schedule with T = 4 epochs and constant LR after epoch 3;
    \item selected bert-base-uncased as best performance / speed ratio;
    \item the best model on validation selection each 0.1 epoch by AUC.
\end{itemize}

For the second step:
\begin{itemize}
    \item hold-out $\approx14\%$ of data to train ensemble of models later;
    \item out-of-5-fold training on the residual $\approx86\%$ of data;
    \item 4 epochs, 512 tokens max length, 16 batch size, 10 gradient accumulation, mixed precision FP16;
    \item default AdamW with lr = 4e-5, Layer-wise Decreasing Layer Rate with decay $\eta = 0.95$ and cosine LR schedule with T = 4 epochs;
    \item the best model on validation selection each 0.1 epoch by F1-score.
\end{itemize}

The final solution contains $N \times K$ models, where $N$ is the number of different backbone BERT architectures, $K$ is the number of folds (5 in the current experiments). Obtained models are further to be ensembled using different strategies with validation on the single hold-out dataset:

\begin{itemize}
    \item hard voting: final spans are selected as at least one model (spans union), as all the models (spans intersection) or as some intermediate methods with at least $m$ models;
    \item soft voting: final probability is calculated as a weighted sum of models probabilities;
    \item train meta classifier.
\end{itemize}

\section{Experiments and Results}
Three pre-trained backbone BERT architectures were considered: BERT base uncased, BERT large uncased \citep{DBLP:journals/corr/abs-1810-04805}, and BERT base pre-trained for Hate Speech Detection \citep{aluru2020deep}. First step setup and results:
    
\begin{itemize}
    \item select subset of Jigsaw toxic classification data: all the targets with toxicity score $\geq$ 0.5 ($L = 135168$ objects) as class 1 and randomly sampled $3 * L$ objects with toxicity score $<$ 0.5 as class 0;
    \item stratified 80\% train, 20\% validation;
    \item 0.968 AUC bert-base, 0.968 AUC bert-large, 0.942 AUC dehate bert.
\end{itemize}

So as models except BERT base uncased did not show compatible performance for token classification (and later for tests on the fold 0 didn't show good F1-score for the actual task as well), later experiments were continued only for BERT base uncased pre-trained model fine-tuned on token classification.

For step two results are following:

\begin{itemize}
    \item train + trial, 8621 comments;
    \item average F1-score over 5 folds is 0.6714.
\end{itemize}

The experiments were conducted with Huggingface transformers library \citep{wolf2019huggingface}.

Many patterns in our results are expected, but some stand out. In general, our model is good at detecting obscene language and utterances that demean honor and dignity or denote low moral character. We noticed that our model is not very good at identifying the posts that have no toxic span annotations. According to the corpus description, in some toxic posts, the core message is conveyed may be inherently toxic. Thus, a sarcastic post can indirectly claim that people of a particular origin are inferior. Hence, it was difficult to attribute the toxicity of those posts to particular spans. In such cases, the corresponding posts were labeled as not containing toxic spans. Among our results, there are many examples where the model detected spans in not annotated posts, for example:

\begin{itemize}
\item ``uhhh Hillary Clinton is a serial killer and thief'': [] (true annotation), [26, 27, 28, 29, 30, 31, 33, 34, 35, 36, 37, 38, 44, 45, 46, 47, 48] (our annotation, ``uhhh Hillary Clinton is a \textbf{serial killer} and \textbf{thief}'');
\item ``This goes way beyond just being an asshole skipper, dude must have some serious mental issues'': [] (true annotation), [35, 36, 37, 38, 39, 40, 41] (our annotation, ``This goes way beyond just being an \textbf{asshole} skipper, dude must have some serious mental issues'').
\end{itemize}

In addition, some texts in the dataset raise questions of the annotation credibility, for example:

\begin{itemize}
    \item ``How the hell is this news?  Am I supposed to be shocked that the Crown Prince of Bahrain or one of the world's biggest celebrity superstars get's better access to the State \underline{Department} then I do?  During which administration has this ever not been true?  
    The media's desperation to keep this election close is far past ridiculous'' (training set, the toxic span annotation is underlined);
    \item ``Yup. NVN u\underline{sed the Pr}ess. The Press was USED. Used like their sister on prom night!
    \underline{Idiots}. All faux-erudite, not realizing they were being played'' (training set, the original annotation is underlined);
    \item ``And you are a complete \textbf{moron} w\underline{ho obv}iously doesn't know the meaning of the word \textbf{narcissist}. By the way your bias is showing'' (test set, the original annotation is underlined, the annotation of our model is highlighted in bold).
\end{itemize}

The final result of our model is presented in Table 1. As can be seen from the table, the systems of the participants produce close results. Our system achieved 67.55\% of F1-score on the test set of this shared task that attracted 91 submitted teams in total. This value exceeded the average result by almost 10\%. 

\begin{table}[]
\caption{Results on the test set.}
\begin{center}
\begin{tabular}{|l|l|l|}
\hline
Rank & Team & F1-score \\ \hline
1 & HITSZ-HLT & 0.7083 \\ \hline
26 & UAntwerp & 0.67552 \\ 
\textbf{27} & \textbf{MIPT-NSU-UTMN} & \textbf{0.67551} \\ 
28 & NLRG & 0.67532 \\ \hline
 & Avg result & 0.57805 \\ \hline
\end{tabular}
\end{center}
\end{table}

\section{Conclusion}

This paper introduces our BERT-based model for toxic spans detection. As expected, pre-training of the BERT model using an additional domain-specific dataset improves further toxic spans detection performance. Experimenting with different fine-tuning approaches has shown that our BERT-based model benefits from the two-step knowledge transfer technique. An ensemble with spans intersection obtained our best result on the test data.
 
In our future work, we will evaluate various language models, such as distilled versions of BERT \citep{sanh2019distilbert,jiao2020tinybert} and RoBERTa \citep{liu2019roberta}.


\bibliographystyle{acl_natbib}
\bibliography{anthology,acl2021}


\end{document}